# Nudging Nutrition: Lessons from the Danish “Fat Tax”

### v2

@@ 4.4 Price mechanism
Since the tax was imposed on producers, it remains unclear to what extent consumers experienced price increases for the taxed products. In this section, we aim to estimate the causal effect of the tax on the prices paid for each product. One limitation of this approach is that we can only observe prices for products that have been purchased, not the overall price level. However, the results presented in Table 5 provide us with insights into whether consumers encountered higher prices for the products they chose to purchase.
@@
Table 5 shows that prices for butter, cream, margarine, and sour cream significantly increased in response to the tax by 23.44%, 10.58% , 11.35%, and 12.31%, respectively. These price increases are lower than those anticipated by the Danish Tax Committee, who anticipated increases in consumer prices up to 35% for animal fat, 33% for dairy, and 24% for margarine. The price increases to a large extent explain our previous finding that consumer expenditure on these products increased by 16%, 22%, and 10%, respectively. The rise in prices corresponded with a notable expenditure increase for both butter and margarine products. Evidently, consumers were willing to pay higher prices for these particular products. The price increase persisted also after the tax was abolished for butter and sour cream, while we find an overall decrease for margarine. We do not find a significant increase in price for other products during the tax period. Interestingly, we observe a substantial decrease in liver sausage prices of about 18%.
@@
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Tax period versus pre-tax quarter Danish fat tax (ATT) -1.918 $14.935^{***}$ $3.610$ $3.968^{***}$3.076 $3.566^{***}$ 1.976 $3.863^{***}$ (se) (2.916) (0.558) (1.660) (0.408) (2.478) (0.509) (3.292) (0.586) % change -2.81% 23.44% 3.92% 10.58% -4.95% 11.35% 1.65% 12.31% obs. 16,863 32,189 27,511 27,735 22,726 28,682 20,925 19,590 p-val. pre-trend 0.554 0.000 0.683 0.009 0.718 0.441 0.833 0.000 Panel B: Post-tax period versus last taxed quarter Danish fat tax (ATT) -3.444 $-10.671^{***}$2.434 $-1.701^{***}$3.917 $-4.054^{***}$3.465 $-1.898^{***}$ (se) (2.894) (0.676) (1.996) (0.474) (2.716) (0.569) (3.595) (0.530) % change -4.82% -15.79% -2.58% -4.18% -6.27% -13.03% -2.92% -5.83% obs. 12,568 22,942 19,930 19,452 16,170 19,230 15,208 13,678 p-val. pre-trend 0.787 0.000 0.489 0.000 0.538 0.032 0.533 0.000 Panel C: Post-tax period versus pre-tax quarter Danish fat tax (ATT) -6.734 $2.039^{**}$ 0.086 -0.093 $-11.091^{***}$ $-1.483^{**}$5.631 $2.585^{***}$ (se) (2.889) (0.583) (1.644) (0.371) (2.414) (0.484) (3.351) (0.572) % change -9.67% 3.22% 0.09% -0.24% -17.95% -5.01% -4.73% 8.15% obs. 19,235 36,058 31,104 30,979 25,609 31,466 23,729 21,739 p-val. pre-trend 0.723 0.000 0.757 0.009 0.615 0.557 0.820 0.000 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
Overall, we do not find a significant increase in prices for products where we observed a short run decrease in consumption. Since we use prices of products purchased by households, the effect on prices might actually be downward biased, because households might decide not to buy when confronted with excessively high prices.
@@ 5 Robustness checks
In this section, we conduct several robustness checks to assess the reliability of the main findings presented in the previous section, addressing the potential concerns with respect to our identification strategy stated in Section 3.2. First, the issue of potential cross-border shopping merits attention, as it might obscure the effect of the tax. To investigate potential cross-border shopping effects, we exclude Southern-Danish households from South/West Jutland and Funen, focusing on Danish households that are less likely to engage in cross-border shopping and hence fully subject to the tax. Interestingly, the results in Table 11 in Appendix D.1 exhibit similar patterns as our main results, indicating that changes in consumption behavior were similar for Danish households that were able to cross-border shop compared to Danish households that did not have this opportunity. Note that in our main results, the unintended effect of cross-border shopping is incorporated as we do include purchases made abroad.
Second, we consider concerns surrounding disparities in Danish and Northern-German household characteristics and assess whether differences in household characteristics between Danish and Northern-German households produced the effects of our main results in Section 4. To gauge the influence of the covariates, we contrast the treatment effects obtained under the conditional parallel trends assumption with those derived under the unconditional parallel trends assumption. If doubly robust DID estimates based on the conditional parallel trends assumption resemble the estimates based on the unconditional parallel trends assumption, this would indicate that covariates contain limited information on consumer behavior trends and are unlikely to drive the results in Section 4. For brevity, we only included the results for the weight outcome, but similar observations apply to the other two main outcomes. Interested readers may obtain the complete set of results upon request. Reassuringly, Figure 39 in Appendix D.2 display extremely similar trends compared to the figures in Appendix C. This suggests that covariates have minimal influence on the trends. Upon comparing the results, it becomes evident that, unsurprisingly, incorporating covariates enhances precision.
@@ 6 Heterogeneity analysis by income levels
Criticisms of the tax being regressive suggest that low-income households may be disproportionately affected compared to wealthier ones. Groceries constitute a larger proportion of the income of poorer households, making them disproportionately affected by the tax if they do not adjust their consumption behavior (Allais et al., , 2010; Zhen et al., , 2011). High-income households, on the other hand, might respond more to health incentives rather than monetary ones. Understanding how the tax affects different income groups is vital for targeted policy improvements, and extrapolating the tax’s applicability to other countries.
@@
In order to investigate potential heterogeneity in the effects of the Danish fat tax, we estimate the effect in two different subsets of the data. First, we estimate the effects of the tax on our three consumption outcomes in the subsample of households who have a low or very low income (refer to Section 3.1 to see how we define income categories). Second, we estimate the same effects but for households who have high or very-high income. The results are presented in Figure 10. Interestingly, Panel (a) of Figure 10 shows a significant reduction in bacon consumption for low-income households, for whom we also observe a significant decrease in consumption of cheese and liver sausage. The effects are not statistically significant for these products, for neither medium-income nor high-income households. Moreover, we find that low-income and medium-income households also reduced their expenditure on cheese and liver sausage, as shown in Panel (b) of Figure 10.
@@
Similar to the main analysis, we observe an uptick in expenditure for both butter and margarine. Notably, the increase is fairly similar between households with high and low incomes, but not statistically significant for medium-income households. This means that low-income households were disproportionately affected by the tax.
@@
The results in Figure 11 show that, after the abolishment of the tax, the increase in butter, cream, margarine, and sour cream consumption is largely driven by high-income households, with relatively weaker effects on low-income households. Interestingly, we find a decrease in the consumption of bacon only for low-income households. However, the effect is marginally significant. [^13] For the interested reader, Figure 40 in Appendix E provides the heterogeneity results for the abolishment effects when comparing the post-tax period to the last taxed quarter.

### v3

@@ 4.5.1 Price mechanism
Since the tax was levied on producers, it remains unclear to what extent consumers experienced price increases for the taxed products. In this section, we aim to estimate the causal effect of the tax on the prices paid for each product. One limitation of this approach is that we can only observe prices for products that have been purchased, not the overall price level. However, the results presented in Table 3 provide us with insights into whether consumers encountered higher prices for the products they chose to purchase.
@@
Table 3 shows that during the tax period, prices for butter, cream, margarine, and sour cream significantly increased in response to the tax by 23.44%, 10.58%, 11.35%, and 12.31%, respectively. These price increases are lower than those anticipated by the Danish Tax Committee, which expected increases in consumer prices up to 35% for animal fat, 33% for dairy, and 24% for margarine.
These price increases largely explain our previous finding that consumer expenditure on these products increased during the tax period. The rise in prices corresponded with notable expenditure increases for butter, cream, margarine, and sour cream, as consumers continued purchasing these products despite higher prices, indicating inelastic demand.
Interestingly, we do not find significant price increases for other products during the tax period. The prices for bacon, cheese, liver sausage, and salami did not change significantly, suggesting that the tax was not fully passed through to consumers for these products, or that market dynamics mitigated the price impact.
After the tax was abolished, as shown in Panel B of Table 3, we find that the price increases persisted for butter and sour cream, with prices remaining 3.22% and 8.15% higher than pre-tax levels, respectively. These increases are statistically significant at the 5% and 1% levels, respectively. In contrast, the price of margarine decreased by 5.01% compared to pre-tax levels (p-value = 0.064), indicating that any price increase during the tax period was reversed after the tax was removed.
Moreover, we observe a substantial decrease in the price of liver sausage after the tax was abolished, with prices dropping by 23.49% compared to pre-tax levels (p-value = 0.005). This suggests that factors other than the tax influenced the price of liver sausage during the post-tax period.
Overall, these findings indicate that the tax led to significant price increases for certain products, which were not entirely reversed after the tax was abolished. The persistence of higher prices for butter and sour cream suggests potential price rigidity or other market factors maintaining higher prices. The absence of significant price changes for other products during the tax period highlights the complexity of tax pass-through in different product markets.
@@
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Tax period versus pre-tax quarter Danish fat tax (ATT) -4.375 $14.935^{***}$ 3.600 $3.968^{***}$2.798 $3.566^{***}$ 0.875 $3.863^{***}$ (se) (3.356) (0.608) (1.697) (0.396) (4.070) (0.484) (3.551) (0.625) p-value 0.362 0.000 0.152 0.000 0.628 0.000 0.868 $0$ change -6.41% 23.44% 3.85% 10.58% -4.50% 11.35% 0.66% 12.31% obs. 15,723 32,189 26,964 27,735 19,521 28,682 19,094 19,590 p-val. pre-trend 0.628 0.000 0.586 0.009 0.892 0.441 0.805 0.000 Panel B: Post-tax period versus pre-tax quarter Danish fat tax (ATT) -7.215 $2.039^{**}$ 0.126 -0.093 $-14.521^{***}$ $-1.483^{*}$ $-9.527^{*}$ $2.585^{***}$ (se) (3.205) (0.633) (1.764) (0.381) (3.558) (0.526) (3.502) (0.577) p-value 0.121 0.031 0.962 0.868 0.005 0.064 0.065 0.002 % change -10.37% 3.22% 0.13% -0.24% -23.49% -5.01% -7.33% 8.15% obs. 18,034 36,058 30,553 30,979 21,942 31,466 21,746 21,739 p-val. pre-trend 0.745 0.000 0.674 0.009 0.964 0.557 0.728 0.000 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
4.5.2 Package sizes
Since the Danish fat tax was levied on producers based on the saturated fat content of their products, it is important to examine not only consumer responses but also potential adjustments made by producers. One plausible producer response to the tax is to reduce package sizes. Smaller package sizes could allow producers to keep the overall price per package lower, even if the price per unit increases due to the tax. This strategy might help prevent a decline in sales volumes by making products appear more affordable to price-sensitive consumers.
In Table 4, we present the causal effects of the Danish fat tax on the average package sizes of various high-fat food products. During the tax period, significant reductions in average package sizes are observed for several products. For bacon, the average package size decreased by 31.46 grams, representing a 16.35% reduction compared to the pre-tax period, with the result being statistically significant at the 5% level (p-value = 0.026). This suggests that producers may have reduced the package sizes of bacon in response to the tax, potentially to keep the total price per package at a level acceptable to consumers.
Similarly, for cheese, there is a significant decrease in average package size by 66.18 grams, an 11.69% reduction (p-value = 0.000). This substantial decrease indicates a strategic adjustment by producers, as the reduction in package size could help maintain sales by making the products appear less expensive, despite the higher per-unit cost.
For liver sausage, the average package size decreased by 20.24 grams, corresponding to a 8.22% reduction, which is statistically significant at the 5% level (p-value = 0.035). This finding aligns with the pattern observed for bacon and cheese, suggesting a consistent producer response across different types of processed meats affected by the tax.
In the case of salami, there is a marginally significant reduction in average package size by 16.05 grams, a 12.26% decrease (p-value = 0.095). While the result is significant at the 10% level, it still points toward a possible producer strategy of offering smaller packages during the tax period.
For other products, such as butter and margarine, the reductions in average package sizes during the tax period are not statistically significant. Butter shows a decrease of 11.77 grams (5.75%, p-value = 0.104), and margarine decreases by 20.68 grams (5.88%, p-value = 0.177). The lack of statistical significance suggests that producers may not have adjusted package sizes for these products in response to the tax, or that the adjustments were not substantial enough to be detected.
After the tax was abolished, some of the reductions in package sizes persisted into the post-tax period. For bacon, the average package size remained significantly lower than in the pre-tax period, with a reduction of 29.79 grams (15.74%, p-value = 0.032). This persistence suggests that the changes made during the tax period may have led to lasting shifts in producer strategies or consumer preferences.
In the case of butter, the average package size decreased significantly by 20.55 grams in the post-tax period, a 10.11% reduction compared to the pre-tax period (p-value = 0.003). This significant decrease, which was not statistically significant during the tax period, indicates that producers adjusted package sizes after the tax was abolished. One possible explanation is that producers delayed adjustments during the tax period due to existing packaging inventories or contracts and implemented changes afterward to optimize production and pricing strategies.
For cheese, the significant reduction in average package size continued into the post-tax period, with a decrease of 58.64 grams (10.64%, p-value = 0.002). The sustained smaller package sizes suggest that both producers and consumers may have adapted to these changes, leading to a new norm in packaging practices for cheese products.
For liver sausage, margarine, salami, cream, and sour cream, the changes in average package sizes during the post-tax period are not statistically significant.
The observed reductions in package sizes, particularly for bacon, cheese, and butter, can be interpreted as a producer response to the fat tax. By offering smaller packages, producers could keep the total price per package lower, making the products more attractive to consumers who are sensitive to price increases. This strategy might help sustain sales volumes despite the higher per-unit costs imposed by the tax. From a consumer perspective, smaller package sizes could lead to reduced consumption if consumers do not compensate by purchasing additional units. However, if consumers adjust by buying multiple smaller packages, the intended effect of the tax on reducing consumption may be diminished.
The persistence of smaller package sizes after the tax was abolished suggests that the changes may have resulted in lasting alterations to the market. Producers might have found that smaller packages are more profitable or better aligned with consumer preferences, or they may have decided not to revert to previous packaging sizes due to the costs associated with changing production processes.
These findings highlight the importance of considering producer behavior when evaluating the impact of taxation policies. Producer strategies, such as adjusting package sizes, can influence how taxes affect consumer prices and purchasing patterns. If producers respond to taxes by modifying product characteristics, the effectiveness of the tax in achieving public health objectives may be affected. Policymakers should take into account potential producer responses when designing tax policies and consider accompanying measures that address these responses to enhance the overall effectiveness of the intervention.
In conclusion, the Danish fat tax appears to have prompted producers to reduce package sizes for certain high-fat products, particularly bacon, cheese, and butter. These adjustments likely aimed to mitigate the impact of the tax on consumers by keeping the total price per package more affordable. The persistence of these changes after the tax was abolished indicates that the tax may have had lasting effects on market practices. Understanding producer responses is crucial for assessing the full impact of taxation policies and for developing comprehensive strategies to promote healthier consumption patterns.
Table 4: The causal effect of the Danish fat tax on the average package sizes.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Tax period versus pre-tax quarter Danish fat tax (ATT) $-31.460^{**}$ $-11.773$ $-66.176^{***}$ 1.176 $-20.241^{**}$ $-20.681$ $-16.048^{*}$ $-5.893$ (se) (9.825) (4.947) (12.638) (6.676) (6.487) (10.221) (7.035) (7.121) p-value 0.026 0.104 0.000 0.906 0.035 0.177 0.095 0.562 % change -16.35% -5.75% -11.69% 0.41% -8.22% -5.88% -12.26% -2.55% obs. 28,874 40,929 32,020 39,274 29,417 38,687 30,800 36,121 p-val. pre-trend 0.713 0.416 0.360 0.065 0.872 0.135 0.846 0.229 Panel B: Post-tax period versus pre-tax quarter Danish fat tax (ATT) $-29.789^{**}$ $-20.545^{***}$ $-58.641^{***}$ 11.990 $-4.845$ $-3.388$ $-9.116$ 0.291 (se) (9.627) (4.717) (13.244) (6.367) (6.123) (9.605) (7.010) (7.175) p-value 0.032 0.003 0.002 0.202 0.606 0.814 0.389 0.978 % change -15.74% -$1$0.64% 4.14% -2.01% -0.98% -7.05% 0.13% obs. 32,648 45,295 35,949 43,648 33,178 42,954 34,711 40,299 p-val. pre-trend 0.468 0.396 0.439 0.038 0.914 0.093 0.857 0.179 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
@@ 5 Robustness checks
In this section, we conduct several robustness checks to assess the reliability of the main findings presented in the previous section, addressing the potential concerns with respect to our identification strategy stated in Section 3.1. First, the issue of potential cross-border shopping merits attention, as it might obscure the effect of the tax. To investigate potential cross-border shopping effects, we exclude Southern-Danish households from South/West Jutland and Funen, focusing on Danish households that are less likely to engage in cross-border shopping and hence fully subject to the tax. Interestingly, the results in Table 15 in Appendix E.1 exhibit similar patterns as our main results, indicating that changes in consumption behavior were similar for Danish households that were able to cross-border shop compared to Danish households that did not have this opportunity. Note that in our main results, the unintended effect of cross-border shopping is incorporated as we do include purchases made abroad.
Second, we consider concerns surrounding disparities in Danish and Northern-German household characteristics and assess whether differences in household characteristics between Danish and Northern-German households produced the effects of our main results in Section 4. To gauge the influence of the covariates, we contrast the treatment effects obtained under the conditional parallel trends assumption with those derived under the unconditional parallel trends assumption. If doubly robust DID estimates based on the conditional parallel trends assumption resemble the estimates based on the unconditional parallel trends assumption, this would indicate that covariates contain limited information on consumer behavior trends and are unlikely to drive the results in Section 4. For brevity, we only included the results for the weight outcome, but similar observations apply to the other two main outcomes. Interested readers may obtain the complete set of results upon request. Reassuringly, Figures 39 in Appendix E.2 and 48 in Appendix E.3 display extremely similar trends compared to the figures in Appendix D. This suggests that covariates have minimal influence on the trends. Upon comparing the results, it becomes evident that, unsurprisingly, incorporating covariates enhances precision.
@@ 6 Heterogeneity analysis by income levels
Criticisms of the tax being regressive suggest that low-income households may be disproportionately affected compared to wealthier ones. Groceries constitute a larger proportion of the income of poorer households, making them disproportionately affected by the tax if they do not adjust their consumption behavior (Allais et al., , 2010; Zhen et al., , 2011). Understanding how the tax affects different income groups is vital for targeted policy improvements, and extrapolating the tax’s applicability to other countries.
@@
In order to investigate potential heterogeneity in the effects of the Danish fat tax, we re-estimate the model for the subsamples of households who have a low or very low income, middle income and high or very high income (refer to Section 3.2 to see how we define income categories). The results are presented in Figure 2. Interestingly, Panel (a) of Figure 2 shows a significant reduction in bacon consumption for low-income households, for whom we also observe a significant decrease in consumption of cheese and liver sausage. The effects are not statistically significant for these products, for neither medium-income nor high-income households. Moreover, we find that low-income and medium-income households also reduced their expenditure on cheese and liver sausage, as shown in Panel (b) of Figure 2.
@@
Similar to the main analysis, we observe an increase in expenditure for both butter and margarine. Notably, the increase is fairly similar between households with high and low incomes, but not statistically significant for medium-income households. This means that low-income households were disproportionately affected by the tax.
@@
The results in Figure 3 show that, after the abolishment of the tax, the increase in butter, cream, margarine, and sour cream consumption is largely driven by high-income households, with relatively weaker effects on low-income households.
7 The effect of the tax on cross-border shopping of butter and liver sausage, geo-spatial effect heterogeneity and spillover effects
The introduction of the Danish fat tax in October 2011 created an economic incentive for Danish consumers, particularly those residing near the German border, to engage in cross-border shopping to circumvent the tax. This section presents descriptive evidence on how the tax influenced cross-border purchasing behavior for two high-fat products: butter and liver sausage. We re-estimate the tax’s impact on these products based on geographic proximity to the German border, evaluating the extent to which cross-border shopping mitigated its intended effects on domestic consumption and expenditure. Additionally, we estimate potential spillover effects on German households living close to the border.
To examine the geo-spatial differences in the tax’s effect and assess these spillover effects, we incorporated information on driving distances to the border. For Danish households, we calculated the driving distances in kilometers from each postal code to the nearest of three German shops located close to the Danish border: (1) Dorfstraße 3, 25927 Aventoft; (2) Alte Zollstraße 44, 24955 Harrislee; and (3) Mummendorfer Weg 7, 23769 Fehmarn. These locations are indicated with blue dots in Figure 4. Figure 4(a) illustrates our division of Denmark into three geographical areas, categorized by their distance from the German border.
Similarly, for German households, we computed the driving distances from the administrative centers of their respective Kreise (administrative districts) to the nearest of the three specified shops. This is used to define border and non-border German regions (See Figure 4(b)). In our data, we can identify purchases made abroad, but we cannot be certain that these are specifically made in Germany. However, since the other neighboring country, Sweden, can only be accessed through passing the bridge or by taking a ferry which involves larger traveling costs, it is highly likely that the vast majority of these purchases are made in Germany. We should acknowledge the higher likelihood of misreporting purchases made abroad. Our descriptive analysis assumes that such misreporting behaviors affect butter and liver sausage similarly to all other products. Moreover, we assume that the introduction of the tax does not directly affect misreporting behaviors.
(a) Denmark.
(b) Northern Germany.
Figure 4: Map with highlighted regions for bordering regions and blue dots are the three German shops used for determining minimum travel distance in km for Danish households. Regions are based on driving distance to the border.
Table 6 presents the fraction of Danes in the sample based on the distance to the closest German shop. Only 2.93% of our sample live within 50 km from the border. Table 6 presents similar overview for Northern-German households.
Table 5: Fraction of Danes in sample based on distance to closest German shop.
Distance Nr. households Fraction
$<$50 km 76 2.97%
50-100 km 232 9.05%
100-150 km 441 17.21%
150-200 km 1124 43.85%
200-250 km 417 16.27%
>=>=> =250 km 273 10.65%
Table 6: Fraction of Germans in sample based on distance of administrative centers within each Kreis to closest German shop at the border.
Distance Nr. households Fraction
$<$50 km 176 14.05%
50-100 km 311 24.82%
100-150 km 311 24.82%
>=>=> =150 km 455 36.31%
7.1 Descriptive Evidence of Cross-Border Purchases
To assess the impact of the fat tax on cross-border shopping, we begin by analyzing the proportion of butter and liver sausage products purchased abroad by Danish households and comparing it to the overall share of all products bought abroad. Figure 5 illustrates the temporal trends in the average share of these products bought abroad by households in bordering regions compared to those in non-bordering regions.
(a) Average share of all products purchased abroad.
(b) Average share of Danes’ butter products purchased abroad.
(c) Average share of Danes’ liver sausage products purchased abroad.
Figure 5: X-axis: Minimum distance to German shop in km. Y-axis: Average consumption five quarters before tax (green), five quarters during tax (red), five quarters after tax (blue).
For butter, there is a noticeable increase in the share of products purchased abroad by households in bordering regions during the tax period (Figure 5(b)). The share rises from approximately 2% before the tax to about 8% during the tax, indicating a significant shift towards cross-border shopping. This elevated level persists even after the tax is abolished, suggesting a lasting change in shopping habits.
A more modest increase is observed for liver sausage purchases (Figure 5(c)). Households near the German border increased their share of liver sausage purchases abroad from approximately 1% to nearly 2% during the tax period. Interestingly, the level of cross-border purchasing seems to have risen to 4% in the post-tax period, implying that this behavior was sustained even after the tax was lifted.
To quantify the relationship between geographic proximity to the German border and cross-border shopping, we employ a random effects model using only Danish households:151515 Since our variable of interest, distance to the border, is constant over time, we rely on a random effects model, as a fixed effects model would not be suitable for time-invariant variables.
$y_{iq}$ $=\alpha_{i}+\beta_{1}dist_{i}+\beta_{2}dist_{i}^{2}+\beta_{3}tax_{q}+\beta_{4}\left(dist_{i}\times tax_{q}\right)+\beta_{5}\left(dist_{i}^{2}\times tax_{q}\right)+\beta_{6}\textbf{X}_{i}+\varepsilon_{iq},$ (4)
where $\textbf{X}_{i}$ contains household characteristics (age of head of household, number of kids under 15, gender of diary keeper, educational attainment (ISCED), income level, and household size), $tax$ is dummy for period where tax was in effect, $i$ index for household, $q$ index for quarter, $dist$ is the minimum distance to one of the three German shops, $y$ is the average share of butter purchased abroad per household per quarter. The model reveals that households residing within approximately 70 kilometers of the German border significantly increased their share of butter purchases abroad during the tax period (Figure 6). We do not find any significant change for liver sausage (Figure 7).
Figure 6: Plot of effect of distance on predicted share of butter purchased abroad before tax and during tax estimated with Equation (4).
Figure 7: Plot of effect of distance on predicted share of liver sausage purchased abroad before tax and during tax estimated with Equation (4).
The increase in cross-border purchases, particularly of butter, highlights a key limitation of the Danish fat tax: the ability of households living near the border to avoid the tax. This raises important questions about the design and implementation of such taxes in smaller countries or regions with accessible, untaxed neighboring markets. Policymakers aiming to use consumption taxes to influence behavior may need to consider geographical context and potential leakage effects, as observed here.
7.2 Geo-spatial effect heterogeneity of the tax
In this section, we examine whether the causal effects of the Danish fat tax varied across different regions of Denmark, particularly focusing on whether households closer to the German border reacted differently compared to those further away. Our analysis did not reveal any significant evidence of heterogeneity in either expenditure or consumption for both products, indicating that the tax’s impact was relatively uniform across regions as shown in Figures 8, 10, 9, and 11.
Figure 8: Map with ATT value per region for expenditure in butter. Values that are significant at 5% level are highlighted with black stripes.
Figure 9: Map with ATT value per region for weight consumed of butter. Values that are significant at 5% level are highlighted with black stripes.
Figure 10: Map with ATT value per region for expenditure in liver sausage. Values that are significant at 5% level are highlighted with black stripes.
Figure 11: Map with ATT value per region for weight consumed of liver sausage. Values that are significant at 5% level are highlighted with black stripes.
7.3 The average treatment effect on the neighbors (ATN): German households
In this section, we examine the potential spillover effects of the Danish fat tax on butter and liver sausage prices in Northern Germany.
We estimate the effect of the Danish fat tax on the average prices paid by German households living close to the Danish border, using households living farther away as a control group. As shown in Table 7, we find no evidence of significant spillover effects from the tax on the prices of butter and liver sausage paid by these German households. This suggests that German shops did not adjust the prices of these products in response to the Danish fat tax.
Table 7: This table shows the ATN for Northern-German households.
Variable Estimate Std. Error P-value
ATT on bordering Germans: butter 0.173 0.799 0.557
ATT on bordering Germans: liver sausage -6.052 5.45 0.890